# AlphaMaze: Enhancing Large Language Models’ Spatial Intelligence via GRPO

### v1

@@ IV-C1 Model Performance on MazeBench
IV-C2 Model Evolution During GRPO
[…]
IV-D Qualitative Results
Qualitative analysis of model outputs revealed notable differences in reasoning behavior across the models. The baseline model often produced nonsensical or incomplete movement sequences, frequently failing to reach the target and exhibiting ”hallucinations” by predicting movements that were not valid within the maze structure. The SFT-only model demonstrated improved coherence and step-by-step progression through the maze, but still occasionally struggled with longer or more complex mazes, sometimes getting stuck in loops or making incorrect turns in later steps of the path.
The SFT+GRPO model exhibited the most sophisticated reasoning behavior. In many instances, we observed emergent chain-of-thought patterns in its output, where the model seemed to explicitly consider wall constraints and spatial relationships at each step before predicting the next move. Furthermore, we observed instances reminiscent of ”aha moments” reported in DeepSeek-R1. For example, in some complex mazes, the SFT+GRPO model initially started down a path, then appeared to ”re-evaluate” its trajectory mid-sequence, correcting course and finding a more efficient or correct path. Error analysis showed that the SFT+GRPO model made fewer invalid moves and was more robust to long-context reasoning challenges compared to the SFT-only model. However, limitations remained, particularly in mazes requiring backtracking or more complex spatial planning beyond the immediate next step.
V Discussion
V-A Analysis of GRPO’s Impact on Visual Maze Reasoning
@@
The qualitative analysis provides further insight into the nature of this improvement. The SFT+GRPO model exhibited more pronounced chain-of-thought reasoning patterns and instances of self-correction, indicating that GRPO is not merely fine-tuning the existing SFT policy, but rather encouraging more sophisticated reasoning processes. The reward function, designed to incentivize correctness, valid movements, and structured output, likely plays a crucial role in shaping this behavior. By rewarding successful navigation and penalizing invalid steps, GRPO encourages the model to learn more deliberate and considered movement strategies.
V-B Comparison with DeepSeek-R1 and RL for Reasoning
@@
In contrast, our two-stage training approach—combining Supervised Fine-Tuning (SFT) to establish foundational step-by-step reasoning with Group Relative Policy Optimization (GRPO) for further refinement—effectively equips the distilled model with robust visual maze-solving skills. Even with only 2000 GRPO steps, the SFT+GRPO model achieves a notable improvement, reaching 93% accuracy and exhibiting clear chain-of-thought behaviors along with self-correction during navigation.

### v2

@@ 4.3.1 Model Performance on MazeBench
4.3.2 Model Evolution During GRPO
Figure 1 displays the MazeBench scores (blue crosses) over GRPO steps along with a linear regression trendline (red dashed line) and its ±$\pm 1$ standard deviation bounds. The steady increase in the trendline indicates that GRPO effectively guides the model towards improved maze-solving policies.
4.4   Qualitative Results
Qualitative analysis of model outputs revealed notable differences in reasoning behavior. The baseline model often produced nonsensical or incomplete movement sequences, frequently failing to reach the target and exhibiting ”hallucinations” by predicting movements invalid within the maze structure. The AlphaMaze-SFT model demonstrated improved coherence and step-by-step progression, but still struggled with longer or more complex mazes, sometimes becoming trapped in loops or making incorrect turns in later stages of the solution path.
In contrast, the AlphaMaze-SFT+GRPO model exhibited the most sophisticated reasoning. In many instances, emergent chain-of-thought patterns were observed, with AlphaMaze (two-stage) appearing to explicitly consider wall constraints and spatial relationships at each step before predicting the next movement. Furthermore, outputs occasionally displayed instances reminiscent of the ”aha moments” reported in prior work on DeepSeek-R1. For example, in some complex mazes, AlphaMaze (two-stage) would initially begin along one path, then appear to ”re-evaluate” its trajectory mid-sequence, correcting its course to find a more efficient or correct solution. Error analysis indicated that AlphaMaze (two-stage) made fewer invalid moves and was more robust to long-context reasoning challenges compared to the AlphaMaze-SFT model. However, limitations remained, particularly in mazes requiring backtracking or complex spatial planning beyond the immediate next step.
5   Discussion
5.1   Analysis of GRPO’s Impact on Visual Maze Reasoning
@@
The qualitative analysis provides further insight into the nature of this improvement. The AlphaMaze-SFT+GRPO model exhibited more pronounced chain-of-thought reasoning patterns and instances of self-correction, indicating that GRPO is not merely fine-tuning the existing SFT policy, but rather encouraging more sophisticated reasoning processes. The reward function, designed to incentivize correctness, valid movements, and structured output, likely plays a crucial role in shaping this behavior. By rewarding successful navigation and penalizing invalid steps, GRPO encourages the model to learn more deliberate and considered movement strategies.
5.2   Comparison with DeepSeek-R1 and RL for Reasoning
@@
In contrast, our two-stage training approach—combining Supervised Fine-Tuning (SFT) to establish foundational step-by-step reasoning with Group Relative Policy Optimization (GRPO) for further refinement—effectively equips the distilled model with robust visual maze-solving skills. Even with only 1600 GRPO steps, the model achieves a notable improvement, reaching 93% accuracy and exhibiting clear chain-of-thought behaviors along with self-correction during navigation.
@@ 6   Conclusion
Appendix A Algorithm
This appendix details the algorithm used to generate the maze reasoning dataset with reset demonstrations. The algorithm processes a base dataset of maze navigation problems and augments it with demonstration of incorrect attempts followed by resets and correct solutions.
Algorithm 1 Maze Reasoning Reset Data Generation - Main Process
0:  Base dataset $D$ containing maze problems with:
1:   - Adjacency list representation of $5\times 5$ maze grid
2:   - Origin and target coordinates
3:   - Correct solution path
3:  Augmented dataset with reset demonstrations
4:  Initialize empty datasets $D_{1}$ and $D_{2}$
5:  for all example $e\in D$ do
6:     Extract adjacency list $A$, origin $O$, target $T$, and path $P$ from $e$
7:     Count walls $W$ around origin $O$
8:     if $W=1$ then
9:        Add $e$ to $D_{1}$
10:        Call ProcessOrder1($e$) {See Algorithm 2}
11:     else if $W=2$ then
12:        Add $e$ to $D_{2}$
13:        Call ProcessOrder2($e$) {See Algorithm 3}
14:     end if
15:  end for
16:  Combine processed examples from $D_{1}$ and $D_{2}$ into the final dataset
Algorithm 2 Order-1 Processing (1 wall at origin)
1:  Procedure ProcessOrder1(example)
2:  $WP\leftarrow\emptyset$Initialize wrong paths set}
3:  for all adjacent node $N$ to origin $O$ do
4:     if $N\notin$ correct path $P$ then
5:        for $n\_steps$ from max_n_steps down to $1$ do
6:           Attempt to extend path from $N$ until a dead end or $n\_steps$ are reached.
7:           if path length = $n\_steps$ or a dead end is reached then
8:              $WP\leftarrow WP\cup\{\text{path}\}$
9:              break
10:           end if
11:        end for
12:     end if
13:  end for
14:  for all path $p\in WP$ do
15:     Generate chain-of-thought steps for path $p$.
16:     Add “Heading in wrong direction” message.
17:     Add RESET marker.
18:  end for
19:  Append original correct solution (path $P$).
20:  Format as conversation pairs.
21:  End Procedure
Algorithm 3 Order-2 Processing (2 walls at origin)
1:  Procedure ProcessOrder2(example)
2:  for $n\_steps$ from max_n_steps down to $1$ do
3:     Generate wrong path $WP$ of length $n\_steps$ starting from $O$.
4:     if a valid path $WP$ is found then
5:        Generate chain-of-thought for $WP$.
6:        if $WP$ ends at a dead end (3 walls) then
7:           Add “Hit a dead end” message.
8:        else
9:           Add “Heading in wrong direction” message.
10:        end if
11:        Add RESET marker.
12:        break
13:     end if
14:  end for
15:  Append original correct solution (path $P$).
16:  Format as conversation pairs.
17:  End Procedure